# Latent Multimodal Reconstruction for Misinformation Detection

### v1

@@ 4.2 Adversarial Prompt Selection
Generating realistic misinformation with an LVLM requires a carefully selected generative prompt $p^{gen}$. However, determining which $p^{gen}$ will produce a dataset $\mathcal{D}^{f}$ that leads to better-performing detection models $\mathsf{M}^{d}$ is not straightforward and would require significant resources for an extensive empirical examination. To address this, we propose Adversarial Prompt Selection, which refines the set of generative prompts by assessing their generated captions $C^{f}$ based on their ability to evade detection by a zero-shot LVLM, ensuring that the generated data presents a meaningful challenge for detection models, rather than being trivially misleading or overly simplistic.
First, we evaluate the LVLM’s ability to perform zero-shot detection of miscaptioned images. We experiment with multiple detection prompts for LLaVa, We experiment with multiple detection prompts for LLaVa, varying their formulation (e.g., word synonyms, rephrasing) and compared their zero-shot performance on a subset of 100 real-world “True vs. MC” samples from VERITE. Based on this process, the detection prompt ($p^{dt}$): “The image is captioned as: [CAPTION]. Is the caption truthful or does it contain falsehoods?” was selected for the next stage.
Next, we evaluate a range of generative prompts against zero-shot LLaVa with $p^{dt}$, experimenting with various formulations similar to those used for the detection prompts. For each generative prompt ($\forall j\in J$), LLaVa generates 1,000 captions from a randomly selected subset of 1,000 truthful image-caption pairs from the NewsCLIPpings training set. Then, the LVLM is tasked with classifying both the truthful and generated captions, and we compute the average detection accuracy for each prompt. This process can be formulated as:
$C^{f}_{j}=\text{LLaVa}(I,C^{t}\mid p^{gen}_{j}),\quad\forall j\in[1,J]$ (2)
$\hat{y}_{j}=\text{LLaVa}(I,C_{j}\mid p^{dt}),\quad C_{j}\in\{C^{t}_{j},C^{f}_{j}\}$ (3)
We reason that generated captions produced by simplistic prompts, which are easily detected, indicate the generation of unrealistic misinformation and are unlikely to contribute to training effective $\mathsf{M}^{d}$ models. For instance, generated captions from prompts such as ‘Create a false caption for the image to propagate misinformation’ and ‘Create fake news based on the image’ were detected with 83.5% and 81.0% accuracy, respectively. Conversely, generative prompts that produce undetectable captions, may fail to produce actual misinformation, instead leaving the original captions unchanged or rephrased. For example, the prompts ‘Provide a caption for the image’ and ‘Maintain the original caption of the image’ resulted in low detection accuracy (42.5% and 50.0%) while having low (7%) or perfect (100%) word overlap with the original captions, indicating a generic or truthful image description rather than misinformation.
Through this process, and after manual inspection, we select four generative prompts, $p^{gen}_{1}$, $p^{gen}_{2}$, $p^{gen}_{3}$, and $p^{gen}_{4}$, with varying adversarial zero-shot detection performance: 57%, 62%, 71%, and 75%. These prompts are then used to generate four distinct sets, $\mathcal{D}^{f}_{1}$, $\mathcal{D}^{f}_{2}$, $\mathcal{D}^{f}_{3}$, and $\mathcal{D}^{f}_{4}$, each corresponding to a different set of false captions related to the images of $\mathcal{D}^{t}$. All four prompts are multimodal, as they take into account both the image and its original caption to generate false captions. For ethical considerations, we refrain from disclosing the specific prompts here, as sharing them could contribute to the automated generation of misinformation; however, they will be available upon request under a research-only license.
@@ 4.3 Generated Data Filtering
Through manual examination, we observe that generative prompts produce more coherent and plausible false captions, often better than NES-based methods in terms of logical coherence and knowledge-based consistency. For instance, as shown in Fig.2, manipulating named entities led to noticeable inconsistencies, such as the incorrect claims of the “Indian parliament” being in “Bangladesh” and “Skeldergate” being in “San Francisco”. These examples underscore the limitations of NES-based methods in preserving basic logical and factual consistency. In contrast, the LVLM demonstrated the ability to generate more creative and plausible misinformation. For instance, it misrepresented a demonstration by reframing its purpose, shifting it from anti-austerity to pro-austerity, and further reinforced the false narrative with claims of “successful implementation” and “widespread approval”. Similarly, it distorted the aftermath of the evacuation following heavy flooding, suggesting that people were “enjoying a leisurely swim in the flooded streets” and “turning a disaster into a fun-filled day”.
However, we also observed that LLaVa can occasionally “ramble”, often rephrasing the original caption while adding redundant or superficial details. For instance, given the original caption “The recent economic boom has enabled new projects such as the Union Trade Centre shopping centre in the heart of Kigali”, LLaVa generated: “The Union Trade Centre shopping centre in Kigali is a prime example of the city’s thriving economy with numerous cars parked outside indicating a bustling shopping scene”; thereby rephrasing the original caption and incorporating minor descriptive elements that do not amount to misinformation.
To address this issue, we apply a post-processing filter that removes samples where the length of the generated caption exceeds a specified character threshold $l\in{\{0,5,10,15,25,50,None\}}$ relative to the original, truthful caption. To maintain a balanced dataset, we remove both the generated pairs ($I^{t}_{i},C^{f}_{i}$) and their corresponding truthful pairs ($I^{t}_{i},C^{t}_{i}$) that exceed this threshold. This filtering process retains 4.5%, 19.1%, 27.8%, 34.9%, 47.9%, 74.0%, and 100% of the dataset for $l\in{\{0,5,10,15,25,50,None\}}$, respectively. We empirically evaluate the impact of this filtering on model performance.
@@ 4.4 Dataset Source and Statistics
We use the NewsCLIPpings ‘Merged/Balanced’ version [14], which has been shown to be effective for OOC detection [22, 24, 21, 43], as the source dataset $\mathcal{D}^{t}$ from which to generate $\mathcal{D}^{f}$, utilizing only its truthful data: 35,536 pairs for training, 3,512 for validation, and 3,512 for testing. After generating $\mathcal{D}^{f}_{1}$, $\mathcal{D}^{f}_{2}$, $\mathcal{D}^{f}_{3}$, and $\mathcal{D}^{f}_{4}$ from $\mathcal{D}^{t}$ using LLaVa and the corresponding generative prompts, we merge $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{f}]$ to address the “True vs. MC” task, or integrate the full NewsCLIPpings dataset $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{x},\mathcal{D}^{f}]$ to address the multi-class task. We preserve the original train/validation/test split of the NewsCLIPpings dataset to prevent any data leakage. This results in a total of 106,605, 10,536, and 10,896 samples for training, validation, and testing, respectively, ensuring a balanced distribution across the three classes.
@@ 5 Latent Multimodal Reconstruction (LAMAR)
Our objective is to develop a reconstruction network $\mathsf{M}^{r}$ that utilizes the embeddings $\mathbf{I},\mathbf{C}$ of image-caption pairs to reconstruct the embedding of the original, truthful caption associated with the image. The reconstructed embeddings $\mathbf{\hat{C}^{t}}$ are then integrated within the final detection model, along with the fused modalities, to aid the detection process. Our rationale is that if the reconstructed embedding $\mathbf{\hat{C}^{t}}$ closely match the input embeddings $\mathbf{C}$, then, the pair $\mathbf{I},\mathbf{C}$ is likely to be truthful. On the other hand, significant discrepancies between $\mathbf{C}$ and $\mathbf{\hat{C}^{t}}$ should suggest potential manipulation of $C$. We hypothesize that this process will provide valuable signals for the detection model.
Since reconstruction networks remain under-explored in the context of MMD, we investigate various alternative strategies, broadly categorized into: (1) end-to-end training, and (2) large-scale pre-training. Additionally, we explore attention, gating, and masking mechanisms to refine the integration of the reconstructed embedding into the Detection network. As illustrated in Figure 3, our pipeline consists of a backbone encoder, followed by modality fusion, a transformer-based reconstruction module, integration of the reconstructed embedding, and a final classification layer. We provide a detailed discussion of these components in the following sections.
Figure 3: End-to-end training of the proposed LAMAR architecture, which utilizes a CLIP ViT L/14 backbone encoder and a Transformer encoder for the reconstruction network with element-wise vector operations for enhanced modality fusion. The reconstructed embedding is then integrated into the detection network through various mechanisms (e.g., gate, mask, attention, or no mechanism) which predicts the final verdict. The reconstruction network is optimized using MSE loss to minimize the difference between reconstructed and ground truth caption embeddings, while the detection network is optimized with cross-entropy (CE) loss.
@@ 5.1 Backbone Encoder
We use CLIP ViT L/14 from OpenCLIP  [^2] https://github.com/mlfoundations/open_clip as the backbone multimodal encoder, $\mathsf{E}_{I}$ and $\mathsf{E}_{C}$, to produce image embeddings $\mathbf{I}\in{R}^{768\times 1}$ and text embeddings $\mathbf{C}\in{R}^{768\times 1}$, which are pre-aligned within a shared embedding space.
@@ 5.2 Modality Fusion
with $\mathbf{F}\in{R}^{5\times 768}$.
@@ 5.3 Detection Network
where $\textbf{W}_{0}\in\mathbb{R}^{768\times 768}$ is a fully connected layer followed by a GELU activation, and $\textbf{W}_{1}\in\mathbb{R}^{n\times 768}$ is the final classification layer, with $n=1$ for binary classification or $n=3$ multi-class classification. Bias terms $b$ are included in the model but omitted here for brevity. Here, the operation ‘Flatten’ refers to converting the concatenated vector $[\mathbf{F};\mathbf{\hat{C}^{t}}]$ into a one-dimensional vector before passing it through the hidden layers. The Detection Network is optimized using binary cross-entropy loss for binary classification and categorical cross-entropy for multi-class classification; denoted as $\mathcal{L}_{d}$.
@@ 5.4 Reconstruction Network
$[\mathbf{t}_{\text{CLS}},\mathbf{t}_{F}]=\mathsf{T}([\mathbf{CLS};\mathbf{F}]),\quad\mathbf{\hat{C}^{t}}=\mathbf{t}_{\text{CLS}}$ (6)
where $CLS$ is a trainable classification token that serves as a global representation of all inputs, and its transformation is defined as the reconstructed caption embedding $\mathbf{\hat{C}^{t}}$. The network is optimized using the MSE loss function, as defined in Eq. 1.
@@ 5.4.1 End-to-end (E2E) training
To jointly optimize the reconstruction network $\mathsf{M}^{r}$ and the detection model $\mathsf{M}^{d}$, we explore end-to-end multi-task training, where the entire LAMAR model is trained using a joint loss function, which combines the reconstruction objective $\mathcal{L}_{r}$ (e.g., MSE) and the detection loss $\mathcal{L}_{d}$ (binary or categorical cross-entropy) as $\mathcal{L}=\mathcal{L}_{d}+\mathcal{L}_{r}$.
This approach enables the detector to utilize both the input representation and the reconstructed embedding, while we explore various mechanisms to dynamically adjust the contribution of the reconstructed embedding based on relevance. Specifically, we explore:
@@
where $\mathbf{g}$ is the gate, $S$ the sigmoid function and $\mathbf{W_{g}}\in\mathbb{R}^{768\times 768}$ while the input to Eq.5 is altered to $[\mathbf{F};\mathbf{\hat{C}^{t}}_{gate}]$.
@@
where $\mathbf{m}$ represents binary mask sampled from a Bernoulli distribution, $\mathbf{W_{m}}\in\mathbb{R}^{768\times 768}$ and the input to Eq.5 is altered to $[\mathbf{F};\mathbf{\hat{C}^{t}}_{mask}]$.
@@
where $\mathsf{W}_{Q},\mathsf{W}_{K},\mathsf{W}_{V}\in\mathbb{R}^{768\times 768}$, and $mean$ denotes average pooling across the first dimension. The input to Eq.5 is altered to $[\mathbf{F};\mathbf{\hat{C}^{t}}_{attend}]$.
@@ 5.4.2 Large-scale Pre-Training (PT)
In addition to end-to-end training, we also investigate a large-scale pre-training approach where the reconstruction network is trained exclusively on truthful captions using a large-scale image-caption dataset, VisualNews [49], comprising 1,259,732 truthful image-caption pairs. We explore two pre-training strategies:

### v3

@@ 4.2 Adversarial Prompt Selection
Generating realistic misinformation with an VLM requires carefully selected generative prompts ($p^{gen}$). However, identifying an “optimal” prompt for a task is non-trivial and costly to evaluate exhaustively. To mitigate this, we leverage Adversarial Prompt Selection, which filters prompts based on how well their generated captions ($C^{f}$) evade detection by a zero-shot VLM, ensuring the generated data presents a meaningful challenge.
First, we assess the VLM’s (LLaVA) zero-shot capability as a ‘Detector’ of miscaptioned images. We conduct a grid search over a candidate pool of prompts varying in wording and granularity. These were evaluated on a balanced calibration set ($N=100$) from the VERITE (“True vs. MC”) benchmark. Through this process, we identified the optimal prompt as: “The image is captioned as: [CAPTION]. Is the caption truthful or does it contain falsehoods?”, which achieved $\approx 70\%$ accuracy. This prompt was selected for subsequently experiments as the adversary detection prompt ($p^{dt}$) against which we evaluate candidate generative prompts ($p^{gen}$). Our analysis revealed that the specific phrasing “contain falsehoods” consistently outperformed synonyms such as “manipulated”, “fake”, or “falsified”. Furthermore, explicitly defining terms, such as stating that false captions “may falsely describe origin, context, and/or meaning of the image” or “contain inconsistencies between named entities (people, dates, locations etc.)” hindered performance, yielding only $\approx 62\%$ accuracy.
Figure 2: Adversarial Prompt Selection: An VLM ‘Manipulator’ generates a false caption using the generative prompt $p_{0}^{gen}$. The VLM ‘Detector’ is then evaluated on the zero-shot classification of both the truthful and generated captions. Intermittent lines indicate the prediction on a single sample, while the overall accuracy is calculated across a balanced set of 2,000 samples. In this specific case, the high overall detection accuracy indicates the generated misinformation is too simplistic; thus $p_{0}^{gen}$ is not selected for the creation of a training dataset.
Next, we assess a range of generative prompts by using the VLM as ‘Manipulator’ to produce 1,000 falsified captions ($C^{f}_{j}$) for 1,000 randomly sampled truthful examples from NewsCLIPpings, as expressed in Eq. 2. Generated captions ($C^{f}_{j}$) are assigned the ground truth label ‘False’. To evaluate the quality of the generated misinformation, we then subject both the generated captions ($C^{f}_{j}$) and the original, truthful captions ($C^{t}$ - labeled ‘True’) to zero-shot classification by the VLM-Detector using the selected detection prompt, as expressed in Eq. 3.
$C^{f}_{j}=\text{VLM-Manipulator}(I,C^{t}\mid p^{gen}_{j}),\quad\forall j\in[1,J]$ (2)
$\hat{y}_{j}=\text{VLM-Detector}(I,C_{j}\mid p^{dt}),\quad C_{j}\in\{C^{t}_{j},C^{f}_{j}\}$ (3)
Prompts that result in high detection accuracy often generate simplistic or unrealistic misinformation. For example, as seen in Fig. 2, the prompt “Create a false caption for the image to propagate misinformation” produced: “UFO sighting: Alien spacecraft descends on Iranian airbase, preparing for an intergalactic invasion” for this specific image, and the VLM-Detector yielded high detection accuracy (83.5%). Conversely, prompts yielding very low detection accuracy could mean that the generated captions are truthful, generic or re-phrased captions, failing to introduce actual falsehoods. For example, the generic prompt “Provide a caption for the image” produced “A lone fighter jet soaring through the clear blue sky, leaving a trail of contrails behind” for the same image, with low overall detection accuracy (42.5%). Therefore, we do not select prompts that produce either easily detectable false captions (due to being unrealistic) or those that fail to introduce falsehoods (due to either being truthful or generic descriptions).
Through this process and manual inspection, we select four generative prompts ($p^{gen}_{1}$$p^{gen}_{4}$) that span a range of adversarial difficulty (57%, 62%, 71%, 75% detection accuracy). These are used to generate four datasets ($\mathcal{D}^{f}_{1}$$\mathcal{D}^{f}_{4}$) of falsified captions conditioned on both the image and its original caption. After identifying the most effective generative prompt ($p^{gen}_{3}$) for LLaVA, we use it for Molmo and Llama as well to keep the computational load of our experiments to reasonable levels while testing cross-model generalizability. For ethical reasons, we withhold exact prompts but make them available upon request under a research-only license.
Figure 3: Examples of truthful and generated captions, alongside false captions created via named entity swaps.
@@ 4.3 Generated Data Filtering
Qualitative analysis (such as in Fig. 3) showed that VLMs can produce more coherent and plausible false captions than NES-based methods, especially in terms of logical and factual consistency. For instance, NES manipulations yielded implausible claims like the ‘Indian parliament’ being in ‘Bangladesh’ or ‘Skeldergate’ in ‘San Francisco’, highlighting the limits of surface-level entity substitution. In contrast, VLMs showed the ability to generate more nuanced and believable misinformation. It might, for instance, reframe an anti-austerity protest as pro-austerity, reinforcing the false narrative with phrases like “successful implementation” and “widespread approval”, or distort the context of flood evacuations by suggesting that people were “enjoying a leisurely swim” and “turning a disaster into a fun-filled day”.
Nevertheless, LLaVA occasionally “rambles”, rephrasing the original caption with superficial elaborations rather than introducing actual misinformation. For instance, from the original caption: “The recent economic boom has enabled new projects such as the Union Trade Centre shopping centre in the heart of Kigali”, LLaVA generated: “The Union Trade Centre shopping centre in Kigali is a prime example of the city’s thriving economy with numerous cars parked outside indicating a bustling shopping scene”, which merely rephrases the original with extraneous details. Moreover, if generation often results in longer captions, this could introduce superficial patterns (“shortcuts”), artifacts, that detection models may learn to exploit, instead of learning deeper features that are associated with content factuality and credibility.
To address this, we apply a post-processing filter that excludes samples where the generated caption exceeds the original by a relative character-length threshold $l\in\{0,5,10,15,25,50,\text{None}\}$, retaining 4.5%, 19.1%, 27.8%, 34.9%, 47.9%, 74.0%, and 100% of the LLaVA-$\mathcal{D}_{3}$ dataset, respectively. We empirically evaluate how filtering affects model performance. To maintain class balance, we removed both the generated pair $(I^{t}_{i},C^{f}_{i})$ and its corresponding truthful pair $(I^{t}_{i},C^{t}_{i})$ when this threshold is exceeded.
@@ 4.4 Dataset Source and Statistics
We use the NewsCLIPpings ‘Merged/Balanced’ dataset [22], which has been shown to be effective for OOC detection [1, 31, 27, 29], as the source dataset $\mathcal{D}^{t}$ from which to generate $\mathcal{D}^{f}$, utilizing only its truthful data: 35,536 pairs for training, 3,512 for validation, and 3,512 for testing. After generating $\mathcal{D}^{f}_{1}$, $\mathcal{D}^{f}_{2}$, $\mathcal{D}^{f}_{3}$, and $\mathcal{D}^{f}_{4}$ from $\mathcal{D}^{t}$ using LLaVA and the corresponding generative prompts, we merge $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{f}]$ to address the “True vs. MC” task, or integrate the full NewsCLIPpings dataset $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{x},\mathcal{D}^{f}]$ to address the multi-class task. We preserve the original train/validation/test split of the NewsCLIPpings dataset to prevent any data leakage. This results in a total of 106,605, 10,536, and 10,896 samples for training, validation, and testing, respectively, ensuring a balanced distribution across the three classes. We use the Molmo and Llama generated datasets in the same manner.
@@ 5 Latent Multimodal Reconstruction (LAMAR)
We propose a reconstruction module $\mathsf{M}^{r}$ that uses image-caption embeddings $(\mathbf{I},\mathbf{C})$ to reconstruct the embedding of the original truthful caption $\hat{\mathbf{C}}^{t}$. The reconstructed embedding serves as an additional signal for the detection model: alignment between $\hat{\mathbf{C}}^{t}$ and the input $\mathbf{C}$ indicates a likely truthful pair, while significant divergence may suggest manipulation. We explore two training strategies—end-to-end and large-scale pre-training—and investigate mechanisms like attention, gating, and masking for integration. As shown in Fig. 4, including a backbone encoder, modality fusion, reconstruction, and detection modules.
Figure 4: End-to-end training of the proposed LAMAR architecture. The input caption $C$ may be either truthful or falsified. A CLIP ViT-L/14 encoder and a Transformer-based reconstruction module, enhanced with element-wise vector operations for modality fusion, outputs the reconstructed embedding $\mathbf{\hat{C}^{t}}$. This is integrated into the detection network via some mechanisms (e.g., gate, mask, attention), which outputs the final verdict. The reconstruction module is trained with MSE loss against the ground-truth $\mathbf{C}^{t}$, while the detection module is trained with cross-entropy (CE).
@@ 5.1 Backbone Encoder
We use CLIP ViT L/14 from OpenCLIP  [^4] https://github.com/mlfoundations/open_clip as the backbone multimodal encoder, $\mathsf{E}_{I}$ and $\mathsf{E}_{C}$, to produce image embeddings $\mathbf{I}\in\mathbb{R}^{768\times 1}$ and text embeddings $\mathbf{C}\in\mathbb{R}^{768\times 1}$, which are pre-aligned within a shared embedding space.
@@ 5.2 Modality Fusion
with $\mathbf{F}\in{R}^{5\times 768}$. Concatenation $\mathbf{I};\mathbf{C}]$ provides the original, isolated features, while the element-wise operations $\mathbf{I}+\mathbf{C}$, $\mathbf{I}-\mathbf{C}$, and $\mathbf{I}*\mathbf{C}$ are included to highlight complementarity, differences, and mutual agreement between the visual and textual embeddings, respectively.
@@ 5.3 Detection Network
where $\textbf{W}_{0}\in\mathbb{R}^{768\times 768}$ is a fully connected layer followed by a GELU activation, and $\textbf{W}_{1}\in\mathbb{R}^{n\times 768}$ is the final classification layer, with $n=1$ for binary classification or $n=3$ multi-class classification. Bias terms $b$ are included in the model but omitted here for brevity. Here, the operation ‘Flatten’ refers to converting the concatenated vector $\mathbf{F};\mathbf{\hat{C}^{t}}]$ into a one-dimensional vector before passing it through the hidden layers. The Detection Network is optimized using binary cross-entropy loss for binary classification and categorical cross-entropy for multi-class classification; denoted as $\mathcal{L}_{d}$.
@@ 5.4 Reconstruction Network
$\mathbf{t}_{\text{CLS}},\mathbf{t}_{F}]=\mathsf{T}([\mathbf{CLS};\mathbf{F}]),\quad\mathbf{\hat{C}^{t}}=\mathbf{t}_{\text{CLS}}$ (6)
where $CLS$ is a trainable classification token that serves as a global representation of all inputs, and its transformation is defined as the reconstructed caption embedding $\mathbf{\hat{C}^{t}}$. The network is optimized using the MSE loss function, as defined in Eq. 1. Both input and target embeddings are L2-normalized, constraining representations to a unit hypersphere, allowing the MSE loss to focus on semantic alignment rather than magnitude variations.
@@ 5.4.1 End-to-end (E2E) training
To jointly optimize the reconstruction network $\mathsf{M}^{r}$ and the detection model $\mathsf{M}^{d}$, we explore end-to-end multi-task training, where the entire LAMAR model is trained using both reconstruction loss $\mathcal{L}_{r}$ (e.g., MSE) and detection loss $\mathcal{L}_{d}$ (binary or categorical cross-entropy) as $\mathcal{L}=\mathcal{L}_{d}+\mathcal{L}_{r}$.
Under this setting, the input caption embedding $\mathbf{C}$ may correspond to either a truthful or a manipulated caption, requiring $\mathsf{M}^{r}$ to learn how to handle both cases appropriately. Since $\mathsf{M}^{r}$ and $\mathsf{M}^{d}$ are trained jointly, the detection model can more dynamically modulate the influence of the reconstructed embedding based on its semantic alignment with the inputs. Specifically, we investigate:
@@
where $\mathbf{g}$ is the gate, $S$ the sigmoid function and $\mathbf{W_{g}}\in\mathbb{R}^{768\times 768}$ while the input to Eq.5 is altered to $\mathbf{F};\mathbf{\hat{C}^{t}}_{gate}]$.
@@
where $\mathbf{m}$ represents binary mask sampled from a Bernoulli distribution, $\mathbf{W_{m}}\in\mathbb{R}^{768\times 768}$ and the input to Eq.5 is altered to $\mathbf{F};\mathbf{\hat{C}^{t}}_{mask}]$.
@@
where $\mathsf{W}_{Q},\mathsf{W}_{K},\mathsf{W}_{V}\in\mathbb{R}^{768\times 768}$, and meanmean denotes average pooling across the first dimension. The input to Eq.5 is altered to $\mathbf{F};\mathbf{\hat{C}^{t}}_{attend}]$.
@@ 5.4.2 Large-scale Pre-Training (PT)
In addition to end-to-end training, we also investigate a large-scale pre-training approach where the reconstruction network is trained exclusively on truthful captions using a large-scale image-caption dataset, VisualNews [20], comprising 1,259,732 truthful image-caption pairs. Unlike end-to-end training, in this setting $\mathsf{M}^{r}$ is trained separately using only perturbed captions $C^{f}$ as input. We explore two pre-training strategies: